OpenStudy (anonymous):

Identify the slope. y = 2.5x – 3.5

OpenStudy (anonymous):

The slope would be 2.5 because Slope Intercept form is Y= mx+b M being the slope and b where it crosses the y-axis

TheSmartOne (thesmartone):

y=mx+b where m is slope and b is y-intercept What here in your equation is m? @gage1127

OpenStudy (anonymous):

y=ax+b so the ans is slop=2.5

TheSmartOne (thesmartone):

Also \(\Huge\cal\color{green}{Welcome~to~Open~Study!!!}\) Since you are new here please read this tutorial for new Open Studiers!! @gage1127 http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8